OpenStudy (haley.fieldss):

can some one simply -3m+3(m+6)

likeabossssssss (likeabossssssss):

what do u think it is ?

OpenStudy (maddielb):

Distribute again, -3m+3m+18, combine like terms 18

KjSaif (kjsaif):

−3m+3(m+6) Distribute: =−3m+(3)(m)+(3)(6) =−3m+3m+18 Combine Like Terms: =−3m+3m+18 =(−3m+3m)+(18) =18

likeabossssssss (likeabossssssss):

no dirct aswer plz

likeabossssssss (likeabossssssss):

jtm

OpenStudy (haley.fieldss):

thanks so much

KjSaif (kjsaif):

Answer: =18